Suppose that a consumer started with an initial endowment of two goods. The table below lists the quantities of each good he starts with, and the prices at which these goods are bought and sold in the market. 



  Endowment Quantity. Market Price

Good 1. 14. 5

Good 2 13 2

What is the market value of this consumer's initial endowment?  


1
Expert's answer
2021-10-15T10:21:45-0400

Market Value of Initial Endowment = "\sum" Market Price "\times" Endowment Quantity

"= (14 \times 5) + (13 \times 2) \\

= 70 + 26 \\

= 96"

The market value of this consumer's initial endowment = $96